14. A 10-kg object experiences the following four forces: a 30-N force to the right a 20-N force to the left; a 40-N force upwards; and a 40-N force downwards. There are no other forces. How much work is done by the net force if the object is moved 5 meters to the right? A) 150 J B) 50J CoJ D)-50J C) 0 J D) -50 J E)-150J 15. How much work must be done to compress a horizontal spring of spring constant 4000-N/m by 0.050 m? B)IOJ C)153 D, OJ E). 10J For problems 16-20, consider the following seenario. A small object of mass m-0.050 kg is tied to a string of length 1.5 m. The object is rotated freely along a vertical circle, as shown. It is set into motion such that at the top of the path, the speed of the object is the minimum needed to continue in a circular path. 16. At which point in the objects path is the kinetic energy of the object the greatest? A) A в) в C) C D) D E) None of these 17. At which point in the objects path is the potential energy of the object the A) A в) в с) с D) D E) None of these 18. At which point in the objects path is the tension in the string the greatest? C) C D) D E) None of these A) A 19. What is the total work done by the tension in the string when the object completes one full rotation? A) 4.6J B)6.9J C) Zero D) It cannot be determined 20. What is the total work done by the force of gravity when the object comple one full rotation? D) It cannot be determined A)4.6J B) 6.9J C) Zero

arrow is going towards a for 16-20
Answer #1

14.

The net force on the object = (30-20) i + (40 - 40)j = 10 i N or 10 N to the right

SInce the object moved by 5 m to the right,

so the work done by the net force on the object = 10*5 = 50 J

Option B is correct.

-----------------------------------------------------

15. The work done to compress the spring: W = Kx2/2 = 4000*(0.050)2/2 = 5 J

The work done = 5 J, Option A is correct.

------------------------------------------------------

16 At the bottom(point C) of the circle, the KE of the object is the greatest. (all energy of the particle is KE at C)

Option C is correct.

-----------------------------

17 At the top(point A) of the circle, the potential energy of the object is the greatest.

Option A is correct.

-----------------------------

18 At the bottom(point C) of the circle, the tension in the string is the greatest. (as the speed at C is the highest)

Option C is correct.

-----------------------------

19 The work done by the tension is zero. As its always perpendicular to the path of the object.

Option C is correct.

-----------------------------

20) The work done by the gravity is zero. As the position of the object is the same after a revolution. Gravity is a conservative force. So the work done by the gravity = 0 J.

Option C is correct.

-----------------------------